# Transient dynamical phase diagram of the spin–boson model

### v1

@@ Appendix B Comparison with analytical estimates at low coupling
$\gamma_{FGR}=\frac{\pi}{2}\alpha\Delta^{s}\omega_{c}^{1-s}e^{-\Delta/\omega_{c}}\propto\alpha\left(\Delta/\omega_{c}\right)^{s},$ (7)
$\frac{\gamma_{FGR}}{\Delta}\left(\frac{\omega_{c}}{\Delta}\right)^{s}=\frac{\pi}{2}\alpha\frac{\omega_{c}}{\Delta}e^{-\Delta/\omega_{c}}\approx 14.21\alpha,$ (8)
which agrees with the scaling in Fig. 6 when $\alpha$ is small. Note that the FGR decay rate is only valid for small $\alpha$.
Figure 6: The damping coefficient $\gamma_{1}$ of the coherent part of the time evolution as function of $\alpha$ for different values of $s$. For small $\alpha$ the damping exhibits the predicted scaling $\gamma_{1}\propto\alpha(\Delta/\omega_{c})^{s}$.
Figure 7 compares the numerically found oscillation frequency $\Omega$ with the analytical prediction from [65]. For a direct comparison, the analytical predictions were obtained by numerically solving Eq. (11) from [65] with our form of spectral density $J(\omega)$ and for the specific parameter values used in the simulation. Our results agree with the analytical estimate at sufficiently weak coupling, but not at strong coupling. In particular, the setup from [65] predicts a smeared-out crossover for all $s<1$, rather than the sharp drop of $\Omega$ that was numerically found at large coupling. We also compare our numerical data with the analytical prediction from [65] as a function of cut-off frequency $\omega_{c}$ for $\alpha=0.05$ and different values of $s$. The agreement is best at small values of $\omega_{c}$ and in the near-Ohmic regime, since $\alpha=0.05$ corresponds to weak coupling for large values of $s$, but to strong coupling for small values of $s$.
Figure 7: Left panel: Oscillation frequency $\Omega$ as function of coupling $\alpha$ for two values of $s$ in the regime that exhibits the frequency-driven coherence/incoherence transition. The symbols represent numerical data (solid lines are guides to the eye) and the dashed lines of the corresponding color are analytical predictions from [65]. The approximate theory and numerics agree at weak coupling. The numerically observed sharp drop to zero frequency at the transition is not predicted by theory. Right panel: Oscillation frequency $\Omega$ as function of the cut-off frequency $\omega_{c}$ at fixed $\alpha$ and for different values of $s$. The symbols represent numerical data and the dashed lines of the corresponding color are predictions based on numerically solving Eq. (11) from [65]. The agreement between the approximate theory and numerics is especially good closer to the Ohmic regime, which corresponds to weaker coupling when $\alpha$ is held fixed.
@@ Appendix C Details on the fitting procedure
The numerical data for $\langle\sigma_{z}(t)\rangle$ obtained with inchworm diagrammatic Quantum Monte Carlo is fitted to the heuristic function given by Eq. (4). As illustrated in Fig. 8, the fits match well the overall shape of the data across the different regimes, but in some cases, the error bars on the input data were scaled in order to ensure acceptable values of $\chi^{2}/$d.o.f. The nonlinear fit function has 7 parameters ($\Omega$, $\phi$, $\gamma_{1}$, $\gamma_{2}$, $a$, $b$, $c$). Although the initial condition requires $\langle\sigma_{z}(0)\rangle=a\cos(\phi)+b+c=1$, strictly enforcing this condition led to worse fits. Instead the error bar on the $t=0$ point was set to have very small (typically between $10^{-4}$ and $10^{-8}$) but nonzero. Because the nonlinear fit has many fit parameters, the $\chi^{2}$-function can have multiple minima and thus the fit result may depend on the choice of the initial parameter guess for the $\chi^{2}$-minimization. The overall error bars on the fit parameters are comprised of the propagated statistical errors of the numerical data, and the systematic error related the variability of the minimization procedure.
Figure 8: Examples of fits to the simulation data $\langle\sigma_{z}(t)\rangle$ for $s=0.5$, $\beta\Delta=20$, $\omega_{c}=10\Delta$. Left panel: $\alpha=0.07$ (coherent and delocalized); Right panel: $\alpha=0.11$ (less coherent and localized). The fits agree qualitatively well with the data.
@@ Appendix D Additional plots of fit parameters
Figure 9 shows a detailed analysis of the transient localization transition for all values of $s$ considered. For values $s\gtrsim 0.5$ the offset curves are significantly less sharp than for smaller values of $s$ and extracting the transition points is more challenging, resulting in a wider confidence interval for the transient phase diagram shown in Fig. 4 of the main text. For the equilibrium QPT the critical exponent $\beta$ defined via $\langle\sigma_{z}\rangle\propto(\alpha-\alpha_{c})^{\beta}$ equals $1/2$ for $s\leq 0.5$. The dynamical data is inconsistent with the equilibrium behavior, indicating a scaling exponent above 1.
Figure 9: Left panel: the offset fit coefficient $c$ as a function of the coupling $\alpha$ for $s\geq 0.6$ (data for $s\lesssim 0.5$ is shown in the main text). While for small $s$ the dynamical localization transition is sharp, for larger values of $s$ the curves are smoother and extracting the transition points is more challenging. The right panel shows the offset values for $s\leq 0.5$ on a log-log scale. The dashed line corresponds to the critical exponent of the equilibrium QPT ($\beta=1/2$) and the dot-dashed line corresponds to an exponent of 2, for comparison. It is clear from the plot that the corresponding scaling of the transient phase diagram is inconsistent with the equilibrium exponent.
Figure 10 shows $\Omega$ plotted on a log-log scale, which indicates the existence of a power law.
Figure 10: Oscillation frequency $\Omega$ as function of coupling $\alpha$ for $s\geq 0.5$ on a log-log scale, which seems to indicate a power law behavior. The extracted powers range between 0.28 for $s=0.5$ and 0.78 for $s=0.9$.

### v2

@@ Appendix B Comparison with analytical estimates at low coupling
$\gamma_{FGR}=\frac{\pi}{2}\alpha\Delta^{s}\omega_{c}^{1-s}e^{-\Delta/\omega_{c}}\propto\alpha\left(\Delta/\omega_{c}\right)^{s},$ (S3)
$\frac{\gamma_{FGR}}{\Delta}\left(\frac{\omega_{c}}{\Delta}\right)^{s}=\frac{\pi}{2}\alpha\frac{\omega_{c}}{\Delta}e^{-\Delta/\omega_{c}}\approx 14.21\alpha,$ (S4)
which agrees with the scaling in Fig. S2 when $\alpha$ is small. Note that the FGR decay rate is only valid for small $\alpha$.
Figure S2: The damping coefficient $\gamma_{1}$ of the coherent part of the time evolution as function of $\alpha$ for different values of $s$. For small $\alpha$ the damping exhibits the predicted scaling $\gamma_{1}\propto\alpha(\Delta/\omega_{c})^{s}$.
Figure S3 compares the numerically found oscillation frequency $\Omega$ with the analytical prediction from [65]. For a direct comparison, the analytical predictions were obtained by numerically solving Eq. (11) from [65] with our form of spectral density $J(\omega)$ and for the specific parameter values used in the simulation. Our results agree with the analytical estimate at sufficiently weak coupling, but not at strong coupling. In particular, the setup from [65] predicts a smeared-out crossover for all $s<1$, rather than the sharp drop of $\Omega$ that was numerically found at large coupling. We also compare our numerical data with the analytical prediction from [65] as a function of cut-off frequency $\omega_{c}$ for $\alpha=0.05$ and different values of $s$. The agreement is best at small values of $\omega_{c}$ and in the near-Ohmic regime, since $\alpha=0.05$ corresponds to weak coupling for large values of $s$, but to strong coupling for small values of $s$.
Figure S3: Left panel: Oscillation frequency $\Omega$ as function of coupling $\alpha$ for two values of $s$ in the regime that exhibits the frequency-driven coherence/incoherence transition. The symbols represent numerical data (solid lines are guides to the eye) and the dashed lines of the corresponding color are analytical predictions from [65]. The approximate theory and numerics agree at weak coupling. The numerically observed sharp drop to zero frequency at the transition is not predicted by theory. Right panel: Oscillation frequency $\Omega$ as function of the cut-off frequency $\omega_{c}$ at fixed $\alpha$ and for different values of $s$. The symbols represent numerical data and the dashed lines of the corresponding color are predictions based on numerically solving Eq. (11) from [65]. The agreement between the approximate theory and numerics is especially good closer to the Ohmic regime, which corresponds to weaker coupling when $\alpha$ is held fixed.
@@ Appendix C Details on the fitting procedure
The numerical data for $\langle\sigma_{z}(t)\rangle$ obtained with inchworm diagrammatic Quantum Monte Carlo is fitted to the heuristic function given by Eq. (4). As illustrated in Fig. S4, the fits match well the overall shape of the data across the different regimes, but in some cases, the error bars on the input data were scaled in order to ensure acceptable values of $\chi^{2}/$d.o.f. The nonlinear fit function has 7 parameters ($\Omega$, $\phi$, $\gamma_{1}$, $\gamma_{2}$, $a$, $b$, $c$). Although the initial condition requires $\langle\sigma_{z}(0)\rangle=a\cos(\phi)+b+c=1$, strictly enforcing this condition led to worse fits. Instead the error bar on the $t=0$ point was set to have very small (typically between $10^{-4}$ and $10^{-8}$) but nonzero. Because the nonlinear fit has many fit parameters, the $\chi^{2}$-function can have multiple minima and thus the fit result may depend on the choice of the initial parameter guess for the $\chi^{2}$-minimization. The overall error bars on the fit parameters are comprised of the propagated statistical errors of the numerical data, and the systematic error related the variability of the minimization procedure.
Figure S4: Examples of fits to the simulation data $\langle\sigma_{z}(t)\rangle$ for $s=0.5$, $\beta\Delta=20$, $\omega_{c}=10\Delta$. Left panel: $\alpha=0.07$ (coherent and delocalized); Right panel: $\alpha=0.11$ (less coherent and localized). The fits agree qualitatively well with the data.
@@ Appendix D Additional plots of fit parameters
Figure S5 shows a detailed analysis of the transient localization transition for all values of $s$ considered. For values $s\gtrsim 0.5$ the offset curves are significantly less sharp than for smaller values of $s$ and extracting the transition points is more challenging, resulting in a wider confidence interval for the transient phase diagram shown in Fig. 4 of the main text. For the equilibrium QPT the critical exponent $\beta$ defined via $\langle\sigma_{z}\rangle\propto(\alpha-\alpha_{c})^{\beta}$ equals $1/2$ for $s\leq 0.5$. The dynamical data is inconsistent with the equilibrium behavior, indicating a scaling exponent above 1.
Figure S5: Left panel: the offset fit coefficient $c$ as a function of the coupling $\alpha$ for $s\geq 0.6$ (data for $s\lesssim 0.5$ is shown in the main text). While for small $s$ the dynamical localization transition is sharp, for larger values of $s$ the curves are smoother and extracting the transition points is more challenging. The right panel shows the offset values for $s\leq 0.5$ on a log-log scale. The dashed line corresponds to the critical exponent of the equilibrium QPT ($\beta=1/2$) and the dot-dashed line corresponds to an exponent of 2, for comparison. It is clear from the plot that the corresponding scaling of the transient phase diagram is inconsistent with the equilibrium exponent.
Figure S6 shows $\Omega$ plotted on a log-log scale, which indicates the existence of a power law.
Figure S6: Oscillation frequency $\Omega$ as function of coupling $\alpha$ for $s\geq 0.5$ on a log-log scale, which seems to indicate a power law behavior. The extracted powers range between 0.28 for $s=0.5$ and 0.78 for $s=0.9$.
Figure S7 shows the time-evolution of $\langle\sigma_{z}(t)\rangle$ for $s=0.2$ and $s=0.8$, for some of the $\alpha$-values shown in Fig. 1 of the main text, but plotted as a set of curves, rather than combined into a density plot. The density plot in the manuscript is generated from many such curves using linear interpolation.
Figure S7: Time-evolution of $\langle\sigma_{z}(t)\rangle$ for $s=0.2$ (left) and $s=0.8$ (right) for select values of $\alpha$. This plot shows some of the data shown in Fig. 1 of the main text, but as a regular plot rather than a density plot. Not all $\alpha$ values shown in the main text are shown here, in order not to overcrowd the plots.
For completeness, Fig. S8 shows the remaining fit coefficients obtained by fitting our data to Eq. (4).
Figure S8: Additional fit parameters: phase $\phi$ (top left), overall damping $\gamma_{2}$ (top right), prefactor $a$ (bottom left), and prefactor $b$ (bottom right).